OpenStudy (anonymous):

Solve D = 1/4 (e – 3f) for e

OpenStudy (anonymous):

\[D = 1/4 (e – 3f) \implies 4D=(e – 3f) \implies 4D+3f=e\]Pretty straightforward, I think.